OpenStudy (anonymous):

(a) sketch the slope field for the differential equation,and (b) use the slope field to sketch the solution that passes through the given point. Y'=-x-2 (-1,1)

OpenStudy (anonymous):

I tried plugging -1 in, but i get -1 not 1. I also tried finding y by integrating and it still doesn't work

OpenStudy (anonymous):

I know the answer, but i dont get how to do it

OpenStudy (anonymous):

brb, i will watch it

OpenStudy (rational):

take your time, its a very good video

OpenStudy (anonymous):

I get how to do them when there is also a y in the equation, then you just plug in points

OpenStudy (rational):

\[y'=-x-2\] first derivative gives the slope of tangent line at a particular point on the curve, yes ?

OpenStudy (anonymous):

yes, then you use that slope to draw the slope field

OpenStudy (anonymous):

so would the slope at (-1,1) be -1?

OpenStudy (rational):

supppose the slope is \(c\), then the isocline is \[c = -x-2\] which is same as \[x=-(c+2)\]

OpenStudy (rational):

which gives a family of vertical lines

OpenStudy (anonymous):

Okay, i understand so far

OpenStudy (rational):

fix slope, \(c = 0\) and you get \[x = -(0+2) \implies x = -2 \] so you draw small line segments with slope=\(0\) along the vertical line \(x=-2\)

OpenStudy (rational):

fix different values for \(c\) and draw as many slope segments as you can

OpenStudy (anonymous):

So instead of plugging in different points, you plug in different slope?

OpenStudy (anonymous):

how would we find the y coordinate though?

OpenStudy (rational):